# Unveiling Inefficiencies in LLM-Generated Code: Toward a Comprehensive Taxonomy

### v1

@@ III-D Step 5: Participants feedback and survey analysis
To evaluate the frequency and relevance of the identified inefficiencies, we conducted a survey targeting software practitioners and researchers using LLMs for code generation. Following prior studies [39, 13], we collected email addresses of GitHub users who collaborated on repositories containing LLM-generated code. To identify relevant repositories, we searched GitHub for repositories where at least one file contained code generated by LLM-based tools or models, including Copilot, ChatGPT, Codex, CodeLlama, Llama, and Llama-2. Specifically, we used the following keyword-based queries: association by/with with Copilot, ChatGPT, Codex, CodeLlama, Llama, Llama-2. This process yielded $411$ repositories. Next, we extracted publicly available email addresses of contributors from these repositories using PyDriller [40]. This resulted in $835$ unique email addresses, of which $51$ were unreachable, leaving us with $784$ successfully delivered emails.  [^6] Date of email mining: November 18, 2024. First, the survey outlined a detailed message about its purpose, scope, and estimated completion time. It was divided into two main parts. The first part includes demographic and open-ended questions regarding LLMs participants had used and related inefficiencies they had encountered based on their experience. The second part evaluated the frequency and relevance of the inefficiencies in LLM-generated code, previously defined in Step 4. For each inefficiency category, we provided a general description, followed by their specific subcategories, each with a description and an example code snippet. Participants were then asked two questions per subcategory, using a 5-point Likert scale: (i) how frequently they encountered the inefficiency reported and (ii) how important they believed it was to address it. Additionally, the survey included an optional question inviting participants to report any inefficiency categories or subcategories not mentioned in the taxonomy based on their experience The survey remained open for $4$ weeks, and we sent a reminder email two weeks after the first call. Once the results for each inefficiency pattern were collected, we processed them to evaluate their frequency and relevance. This was achieved by aggregating the results using weighted averages of Likert scale values, following methodologies from similar studies [13]. Additionally, we complemented the analysis with comments from responders to validate the taxonomy.
@@ IV Results
In this section, we present the results of the two RQs investigated in this study. First, we present our taxonomy of LLM-generated code inefficiencies, and then, we discuss the perceptions of our proposed taxonomy through our survey.
@@ IV-A1 Taxonomy
We organized our observations as a taxonomy of five categories with 19 subcategories. In the following, we explain each of the five categories with their respective subcategories. Examples of each can be found in our replication package (code snippets).
A. General Logic. This category encompasses inefficiencies related to the overall logic of a given task, including issues with understanding requirements and translating them into code.
  • Wrong Logic. It occurs when the generated code lacks meaningful logic, containing only comments, hard-coded values, returning method inputs, or using a simple pass statement. It also includes case-by-case implementations that fail to follow coherent logic. Additionally, it encompasses situations where all requirements are implemented incorrectly, with none being addressed properly.
  • Partially Wrong Logic Different from Wrong Logic, this subcategory occurs when some requirements are implemented correctly while others are either incorrectly implemented or not implemented at all.
  • Wrong Method Input. It occurs when the LLM generates a statement that is generally coherent with the logic but uses incorrect or inappropriate inputs. For example, passing a single element to the sorted method instead of a list may lead to unintended behavior.
  • Exception/Corner Case Handling. This occurs when the generated code lacks exception handling or fails to address corner or edge cases. Initially, this sub-category was merged with Partially Wrong. However, following survey participants’ suggestions, it was detached as a standalone subcategory.
B. Performance. This category groups inefficiencies that negatively impact the generated code’s performance.
  • A More Memory-Optimal Solution Exists (Memory). This subcategory captures cases where the same logic can be implemented in a more memory-efficient manner. For example, recursion may be avoided when intermediate results are stored on the call stack, potentially leading to memory overflow for deep recursion. Similarly, list comprehensions, which store all results in memory, can be replaced with generator expressions to enable lazy evaluation and reduce memory usage.
  • A more optimal solution exists (Time). It identifies the cases where an equivalent, more time-efficient algorithm exists compared to the LLM-generated code. In this subcategory, we report four additional subcategories:
    • Sub-Optimal Solution (Time Complexity): It occurs when the implemented logic can be replaced with another algorithm with lower time complexity. For instance, using Quick Sort, with a time complexity of $O(n\log n)$ on average, is generally more efficient than Insertion Sort, which has a time complexity of $O(n^{2})$ in the average and worst cases.
@@
    • Redundant Steps. This occurs when an instruction essential to the logic is unnecessarily repeated. To avoid such cases, the given instruction could be executed once, and its output could be stored in a local variable instead of re-executing the same instruction.
@@
      + Iterative Block (Initialization Condition). It occurs when the initialization condition of a loop is overly broad and can be narrowed without affecting the method’s correctness. For example, when determining whether a number is prime, we can stop checking divisibility at $\sqrt{n}$ rather than iterating all the way up to $n$.
@@
      + Iterative Block (Stopping Condition). It occurs when the stopping condition for a repetitive block is too broad and can be optimized without changing the intended functionality or when it lacks an early stopping condition when applicable. For instance, when searching for an element in a sorted list, instead of iterating through the entire list, we can stop early if the current element exceeds the target value.
@@
C. Readability This category encompasses issues on the overall readability, clarity, and ease of comprehensibility of LLM-generated code snippets.
@@
  • Confusing Variable Naming. It refers to variable names that can confuse, even if they do not result in runtime errors. It includes issues such as redefining existing names (e.g., using dict as a variable or method name) or variable shadowing, where the same variable name is used in different scopes.
@@
  • A more readable method exists. This refers to cases where the code is difficult to understand or does not adhere to conventions. For example, the combination of list comprehension, map, and lambda functions, along with a for loop and multiple method calls, such as the use of this statement: return reduce(lambda x, y: x * y, [int(i) for i in str(n) if int(i) % 2 == 1]).
@@
D. Maintainability. This category highlights inefficiencies in the LLM-generated codes that may negatively impact the maintainability of the code, possibly being intended to be integrated into larger systems.
@@
  • Code Duplication. This occurs when identical or similar blocks of code appear in multiple places without adding new value to the main logic of the code, essentially resembling code cloning [41].
@@
  • Comment Duplication. It occurs when repetitive comments are used excessively, identical or similar, without providing additional insight or explanations about the logic.
@@
  • Conditional Block. It groups inefficiencies related to conditional blocks, i.e., if blocks. Although it does not impact the functionality, it is important to address it for better code clarity and easier testing/debugging [42]. While in our studied sample we did not observe foreach blocks, this subcategory may be extended to cover them as well.
@@
    • Unnecessary Else. It takes place when an else statement is used unnecessarily within an if block and can be removed without affecting the program’s behavior. It usually happens when the preceding if or elif statements already contain a return or break, making the else unnecessary.
@@
    • Unnecessary Conditional Block. It refers to cases where using a conditional block adds unnecessary complexity and can be simplified by directly using the output of a given condition in a return statement.
@@
E. Errors Inefficiencies falling under Erros are related to the fundamental correctness of the code regardless of the logic being implemented. These issues can lead to runtime errors, unexpected behavior, or failure to execute the code properly.
@@
  • Missing Variable Declaration. It occurs when a variable is used without being properly declared, initialized, or passed as a parameter for a method, which may result in runtime errors.
@@
IV-A2 Code Inefficiencies Patterns
TABLE I: (1) Frequency of inefficiencies in LLM-generated code, and (2) Survey-based taxonomy validation Results
@@ IV-A2 Code Inefficiencies Patterns
As previously discussed, our taxonomy of inefficiencies was derived from observed patterns in our studied sample using multi-labeling at both the category and subcategory levels. Regarding category frequency, most samples were labeled with a single category (56.5%), followed by those with two (24.59%) and none (9.96%). A similar trend was observed at the subcategory level, where the majority had one subcategory (54.07%), followed by two (18.29%) and three (11.38%). Notably, 33.54% of the samples exhibited characteristics spanning multiple categories, underscoring the complexity and overlap of inefficiency patterns across different code quality aspects. Table I presents the frequency of inefficiency categories and subcategories (2nd-5th columns). [^7] Since multi-labeling is allowed, the total may exceed 100% General Logic emerges as the most frequent category across all models (68.5% of samples). Among the selected models, CodeGemma, the least efficient, based on Top@k (leaderboard [33]), exhibited the highest rate of logic-related issues (74.39%), often producing code with no meaningful logic (57.93% of its cases). While CodeLlama, the most efficient, based on Top@K, showed fewer critical logic errors (60.37%), though it frequently generated partially correct solutions (47.56%), suggesting even advanced models may misinterpret requirements [43].
Performance inefficiencies ranked second (34.15%) with CodeLlama leading (48.78%). Sub-Optimal (Memory), Sub-Optimal (Time Complexity) and Redundant Steps were prevalent highlighting that even the most efficient model are prone to performance issues such as GPT-4 [44].
Regarding Maintainability issues varied significantly across models. DeepSeek-Coder, trained exclusively on code, exhibited the lowest maintainability flaws (12.8%), whereas CodeLlama and CodeGemma (trained on mixed code/natural language data) showed higher rates (29.88% and 20.73%, respectively). This discrepancy suggests that code-specific training may improve structural quality, while hybrid training may dilute maintainability focus. The prevalence of Unnecessary Else blocks (21.95% in CodeLlama) underscores a broader challenge: LLMs often generate verbose control flows, complicating long-term maintenance [45].
Readability (4.67%) and Errors (5.69%) were less frequent but non-trivial. Syntax errors were rare ($<2.44\%$), confirming LLMs’ strong syntactic grasp. However, Missing Imports (6.1% in CodeLlama) reveal persistent gaps in dependency resolution—a known limitation even in most-efficient models, GPT-4, [46].
@@
IV-A3 Analysis of Co-occurrence Categories
To gain deeper insights into inefficiency relationships, we analyze how different categories co-occur across models. Figure 2 presents the normalized category co-occurrence heatmap across all models. The co-occurrence values are normalized row-wise, with each value divided by the frequency of the corresponding category. We observe strong inter-dependencies between inefficiencies in LLM-generated code. General Logic inefficiencies exhibit the highest co-occurrence rates with other categories, particularly Readability (0.74), Errors (0.64), and Maintainability (0.58). This suggests that flawed logic not only impacts correctness but also contributes to inefficiencies in code clarity, structure, and long-term maintainability. Similarly, Performance inefficiencies frequently co-occur with Readability (0.48) and Maintainability (0.46), highlighting that performance issues are often tied to code complexity and structure, rather than isolated algorithmic inefficiencies. Interestingly, Errors demonstrate weaker co-occurrences with other inefficiencies, with their strongest link being to Maintainability (0.36) indicating that while syntax and dependency errors are less common, they may still hinder maintainability efforts by requiring additional debugging and refactoring. Meanwhile, Readability and Maintainability are closely linked (0.35), reinforcing the idea that clearer code structures contribute to better long-term maintainability. These findings suggest that improving LLM-generated logic and performance could lead to broader improvements across multiple inefficiency categories, reducing the need for manual corrections and enhancing code usability in real-world development.
@@ IV-B RQ2: Relevance of Identified Inefficiencies for Practitioners and Researchers
We kept the survey open for four weeks reaching $58$ responses, resulting in a response rate of $7.39$%, which aligns with typical rates observed in software engineering surveys [26, 47, 48, 49]. Here, we first present the demographic information about the participants and their experience with coding using LLMs. Second, we provide a detailed analysis of the reported frequency and perceived relevance of inefficiencies in LLM-generated code.
@@ IV-B1 Demographic Information and Experience with Coding using LLMs
In our survey, all participants answered the demographic questions, obtaining the following answers.  [^8] Some percentages may exceed 100% as participants were able to select multiple options for some questions Our sample consisted of participants from both academia (60%) and industry (40%). Among academic participants, we had graduate, undergraduate students, and postdoctoral researchers/lecturers (29, 3, and 3, respectively). From the industry, we had 11 software developers, 7 ML/data science developers, 4 senior executive roles, and 1 Cybersecurity developer. About the experience, 37 participants had more than 5 years of experience, while 11 had between 3-5 years, and 10 had between 1-3 years. Regarding adopting LLMs to assist in coding, 98% of the participants report using them. Both proprietary and open-source LLMs are in use, with 68% relying exclusively on proprietary models. Among proprietary models, ChatGPT and Copilot are the most commonly used (89.66% and 29.31%, respectively). On the other hand, for open-source models, the most popular are Mistral, the Phi model family, and Qwen-Coder (5.17%, 3.45%, and 3.45%, respectively). Finally, participants also use a variety of programming languages, with the top five being Python, C/C++, JavaScript, Java, and Go (91.38%, 18.97%, 18.97%, 8.62%, and 3.45% respectively).
Before getting into the questions addressing our taxonomy, we asked participants about their general experience and faced issues in LLM-generated code. They reported inefficiencies that can be categorized into two main groups: user experience (outside the scope of our taxonomy) and code-related, which are directly relevant to our research. We believe such a diverse distribution across academia and industry, combined with a strong representation of experienced professionals, provides a well-balanced perspective on inefficiencies in LLM-generated code. The diversity in roles and expertise levels enhances the reliability of our findings by capturing insights from both research-oriented and real-world coding environments. We discuss the reported inefficiencies as follows.
@@
These inefficiencies focus on the challenges/hardness users face when interacting with LLMs, rather than the generated code. Overall, 6 participants reported that LLMs often perform poorly when tasked with debugging or modifying specific parts of the code, as also reported in previous studies [50]. In the same way, programming language and library familiarity also play a significant role in LLM efficiency. 5 Participants observed that LLMs tend to perform better with popular languages and widely used libraries, such as Python and NumPy, respectively, compared to less popular ones, like DHL and Tinkter [51]. Moreover, the programming language paradigm also matters; one participant noted that LLMs perform better with interpreted languages than with compiled ones. Participants also shared different experiences when using Chat-based LLMs, such as ChatGPT, compared to code assistants like Copilot. One participant mentioned preferring ChatGPT, reporting that Copilot made “stupid suggestions and distracted me”. Conversely, another participant found Copilot more useful for providing inline suggestions while maintaining overall visibility of the codebase findings [52].
@@
Participants reported various inefficiencies in the generated code that align with the focus of our taxonomy. 19 participants stated that LLMs often struggle to produce correct logic on the first attempt, even for simple tasks accompanied by detailed descriptions, requiring multiple iterations to achieve the desired outcome. P22 specifically noted LLMs may generate “unnecessarily long or complex (for simple tasks)” However, for complex tasks, as reported by P15, such as building a Flask application, LLMs tend to“oversimplification to the point of uselessness”, offering high-level steps or generalized code that lack specificity, performance efficiency, and functionality. Furthermore, 3 participants reported that handling exceptions and corner cases is often overlooked. All these reported issues fall under General Logic category of our taxonomy and were introduced spontaneously before participants were introduced to our taxonomy.
Another notable issue is the misalignment between LLM advancements and technological progress. 13 participants observed that LLMs often fail to account for updates in library versions, resulting in version tracking challenges. Examples include mixing methods from different library versions, using deprecated libraries, calling non-existent methods, or generating code incompatible with the specified library or API versions[46]. Some of these issues align with Wrong Method Input (General Logic) and Missing Import (Errors) sub-categories of our taxonomy. Participants—using various programming languages, like Python, C/C++, Java, JavaScript, Go, and Bash—also reported readability-related inefficiencies. They included poor code formatting, failure to adhere to user-defined coding styles or naming conventions, inconsistent or missing variable names, unused variables, and reliance on outdated coding practices. Such observations highlight that these issues may persist across different programming languages [16]. Additionally, 6 participants highlighted performance concerns, while others noticed maintainability issues, noting that the generated code is not always performant and often lacks robust, modular, or object-oriented architecture. Insufficient documentation and comments further exacerbate the maintainability challenges. These observations are reflected in the proposed Readability, Performance, and Maintainability categories of our taxonomy and aligned with findings in the literature [16, 14, 16]. 7 participants noted that context limitations were a significant factor in the issues they encountered. Their experiences with memorization in chat-based LLMs varied. Two participants found that LLMs had a limited context window and frequently forgot past interactions [53]. In contrast, another participant reported that LLMs tended to retain outdated context, stating that they needed to “start a new chat to get rid of hallucinations.”

### v2

@@ III-D Step 5: Participants’ feedback and survey analysis
To evaluate the relevance and gather insights into the perceived popularity of the identified inefficiencies in real-world settings, we conducted a survey targeting software practitioners and researchers using LLMs in their code generation tasks. Following prior studies [43, 13], we collected the email addresses of GitHub users who collaborated on repositories containing LLM-generated code. To identify relevant repositories, we searched GitHub for repositories where at least one file contained code generated by LLM-based tools or models, including Copilot, ChatGPT, Codex, CodeLlama, Llama, and Llama-2. Specifically, we used the following keyword-based queries: association between by/with and Copilot, ChatGPT, Codex, CodeLlama, Llama, Llama-2. This process yielded $411$ repositories. Next, we extracted publicly available email addresses of contributors from these repositories using PyDriller [44]. This resulted in $835$ unique email addresses, of which $51$ were unreachable, leaving us with $784$ successfully delivered emails (mined on November 18, 2024).
To ensure clear communication and maximize participant engagement, we structured the survey to clearly outline its objectives, structure, and expectations. First, the survey outlined a detailed message about its purpose, scope, and estimated completion time. The survey questionnaire is divided into two main parts. The first part consists of demographic and open-ended questions regarding the LLMs participants have used, as well as the inefficiencies they encountered based on their experiences. The second part assessed the frequency and relevance of inefficiencies in LLM-generated code, as defined in Step 4. For each inefficiency category, we provided a general description, followed by its specific subcategories, each accompanied by a description and an example code snippet. Participants were then asked two questions per subcategory, rated on a 5-point Likert scale: (i) how frequently they encountered the inefficiency reported and (ii) how important they believed it was to address it. Additionally, the survey included an optional question, allowing participants to report any inefficiency categories or subcategories not covered in the taxonomy, based on their experience. The survey remained open for $4$ weeks, and we sent a reminder email two weeks after the first call. Once the results for each inefficiency pattern were collected, we processed the results to evaluate their popularity and relevance. This was achieved by aggregating the results using weighted averages of Likert scale values, following methodologies from similar studies [13]. Additionally, we analyzed participants’ feedback to gain deeper insights into their experiences with LLM-generated code. Their comments provided context for interpreting the results and helped identify potential gaps or overlooked aspects in the taxonomy.
@@ IV Results
In this section, we present the findings for our two RQs explored in this study. We first introduce our taxonomy of inefficiencies in LLM-generated code, followed by an analysis of the perceptions of practitioners and researchers on the proposed taxonomy, as gathered through our online survey.
@@ IV-A1 Taxonomy
We organized our observations as a taxonomy of five categories with 19 subcategories. The complete taxonomy is shown in Figure 2. In the following, we explain each of the five categories with their respective subcategories. Examples of each inefficiency subcategory (i.e., code snippets) can be found in our replication package.
Figure 2: LLM-generated Code Inefficiencies: A comprehensive Taxonomy
A. General Logic. This category encompasses inefficiencies related to the general logic of a given task, including inefficiencies in understanding requirements and translating them into code.
  • Wrong Logic. It occurs when the generated code lacks meaningful logic, containing only comments, hard-coded values, returning method inputs, or using a simple pass statement. It also includes case-by-case implementations that do not follow a coherent logic. Additionally, it encompasses situations where all requirements are implemented incorrectly, with none being addressed properly.
  • Partially Wrong Logic Different from Wrong Logic, this subcategory occurs when some requirements requested in the prompt are implemented correctly while others are either implemented incorrectly or not implemented at all.
  • Wrong Method Input. It occurs when the LLM generates a statement that is generally coherent with the logic but uses incorrect or inappropriate inputs. For example, passing a single element to the sorted method instead of a list may lead to an unintended behavior.
  • Exception/Corner Case Handling. This occurs when the generated code lacks exception handling or fails to address corner cases. Initially, this subcategory was merged with Partially Wrong. However, following the suggestions of the survey participants, it was detached as a standalone subcategory.
B. Performance. This category encompasses inefficiencies that degrade the efficiency of the generated code, affecting aspects such as memory management and execution time.
  • Sub-Optimal Solution Exists (Memory) This subcategory captures cases where the same logic can be implemented in a more memory-efficient way. For example, recursion may be avoided when intermediate results are stored on the call stack, potentially leading to memory overflow for deep recursions. Similarly, list comprehensions, which store all results in memory, can be replaced with generator expressions to enable lazy evaluation and reduce memory usage.
  • Sub-Optimal Solution Exists (Time) It identifies the cases where there is an equivalent, more time-efficient algorithm compared to the LLM-generated one. In this subcategory, we report four additional subcategories:
    • Sub-Optimal Solution (Time Complexity): It occurs when the implemented logic can be replaced with another algorithm with lower time complexity. For instance, using Quick Sort, with a time complexity of $O(n\log n)$ on average, is generally more efficient than Insertion Sort, which has a time complexity of $O(n^{2})$ in the average and worst cases —particularly when sorting large lists.
@@
    • Redundant Steps. This occurs when an essential instruction to the logic is repeated unnecessarily. To avoid such cases, the given instruction could be executed once, and its output could be stored in a local variable instead of re-executing the same instruction.
@@
      + Iterative Block (Initialization Condition). It occurs when the initialization condition of a loop is too broad and can be narrowed without affecting the correctness of the method. For example, when determining whether a number is prime, we can stop checking the divisibility at $\sqrt{n}$ rather than iterating all the way up to $n$.
@@
      + Iterative Block (Stopping Condition). It occurs when the stopping condition for a repetitive block is very broad and can be optimized without changing the intended functionality or when it lacks an early stopping condition when applicable. For instance, when searching for an element in a sorted ascending list, instead of iterating through the entire list, we can stop early if the current element exceeds the target value.
@@
C. Readability This category encompasses inefficiencies related to the overall readability, clarity, and ease of comprehensibility of LLM-generated code snippets.
@@
  • Confusing Variable Naming. It refers to variable names that may confuse the developer, even if they do not result in runtime errors. It includes inefficiencies such as using existing identifiers to define new variables (e.g., using dict as a variable or method name) or variable shadowing, where the same variable name is used in different scopes.
@@
  • Sub-readable method exists. This refers to cases where the code is difficult to understand or does not adhere to conventions. For example, the combination of list comprehension, map, and lambda functions, along with a for loop and multiple method calls, such as the use of this statement: return reduce(lambda x, y: x * y, [int(i) for i in str(n) if int(i) % 2 == 1]).
@@
D. Maintainability. This category highlights inefficiencies in LLM-generated code that may negatively impact the maintainability of the code, possibly intended to be integrated into larger systems.
@@
  • Code Duplication. This occurs when identical or similar blocks of code appear in multiple places without contributing to the main logic, essentially resembling code cloning [45].
@@
  • Comment Duplication. It occurs when repetitive comments are used excessively, identical, or similar, without providing additional insight or explanations about the logic.
@@
  • Conditional Block. It groups inefficiencies related to conditional blocks, i.e., if blocks. Although it does not affect the functionality, it is important to be addressed for better code clarity and easier testing/debugging [46]. While in our studied sample we did not observe foreach blocks, this subcategory may also be extended to cover them.
@@
    • Unnecessary Else. It takes place when an else statement is used unnecessarily within an if block and can be removed without affecting the logic. It usually happens when the preceding if or elif statements already contain a return or break, making the else unnecessary.
@@
    • Unnecessary Conditional Block. It refers to cases where using a conditional block adds unnecessary complexity, and the code can be simplified by directly using the output of a given condition in a return statement.
@@
E. Errors Inefficiencies falling under Erros are related to the fundamental correctness of the code regardless of the logic being implemented, like runtime errors, unexpected behavior, or failure to execute the code properly.
@@
  • Missing Variable Declaration. It occurs when a variable is used without being properly declared, initialized, or passed as an argument for a method, which may result in runtime errors.
@@
Overall, LLM-generated code exhibits similar inefficiencies to human-written code. While LLMs often pass more test cases than human developers, they struggle with tasks requiring domain expertise [17]. Regarding performance, although LLMs can generate efficient solutions for certain tasks, they often under-perform compared to human-optimized implementations [47]. In terms of readability, LLM-generated code is generally comparable to human-written code but may deviate from coding standards [16, 48]. However, its increased complexity can elevate maintenance effort [17]. While LLMs can address maintainability issues, their fixes frequently introduce new errors at a higher rate than human corrections [49].
IV-A2 Analysis of Inefficiency Patterns
TABLE I: (1) Frequency of inefficiencies in the studied sample of LLM-generated code, and (2) Survey-based taxonomy validation Results
@@ IV-A2 Analysis of Inefficiency Patterns
Most samples were labeled with a single category (56.5%), followed by those with two (24.59%) and 9.96% were not labeled, as raters found no inefficiencies and considered them efficient. A similar distribution appeared at the subcategory level with 54.07% of samples assigned one subcategory followed by 18.29% with two and 11.38% with three. Notably, 33.54% of the samples spanned multiple categories, highlighting the interconnected nature of inefficiency patterns across different code quality aspects.
Table I presents the frequency of inefficiency categories and subcategories (since multi-labeling is allowed, totals may exceed 100%). General Logic emerges as the most frequent category across all models (68.5% of samples). Among the selected models, CodeGemma, the least efficient, based on Top@k (leaderboard [30]), exhibited the highest rate of logic-related inefficiencies (74.39%), often generating code without a meaningful logic (57.93%). While the most efficient model based on Top@K, CodeLlama, showed fewer critical logic errors (60.37%), though it also frequently generated partially correct solutions (47.56%), suggesting that even advanced models may misinterpret requirements, which is aligned with findings of other research works [50].
Performance inefficiencies ranked second (34.15%), with CodeLlama exhibiting the highest frequency within this category. Sub-Optimal (Memory) (18.7%), Sub-Optimal (Time Complexity) (18.5%) and Redundant Steps (5.28%) were the most prevalent performance-related inefficiencies. These inefficiencies are not exclusive to smaller or open-source models; prior studies have reported similar performance issues in GPT-4-generated code [51], reinforcing the broader challenge of optimizing LLM-generated code across different model architectures.
Maintainability inefficiencies varied significantly between models. DeepSeek-Coder, trained exclusively on code, exhibited the lowest Maintainability issues (12.8%), whereas CodeLlama and CodeGemma (trained on mixed code/natural language data) have higher rates (29.88% and 20.73%, respectively). This discrepancy suggests that code-specific training may improve structural quality, while incorporating natural language data in training may dilute the focus on maintainability. The prevalence of Unnecessary Else blocks (21.95% in CodeLlama) underscores a broader challenge: LLMs often generate unnecessarily complex control flows, which complicate long-term maintenance [52].
Readability (4.67%) and Errors (5.69%) were less frequent but non-trivial. Syntax Errors were rare ($<2.44\%$), confirming the strong syntactic grasp of code LLMs. However, Missing Imports (6.1% in CodeLlama) reveal persistent gaps in dependency resolution, a known limitation even in highly-efficient models like GPT-4 [53].
@@
To gain deeper insights into inefficiency relationships, we analyze how different categories co-occur across models. Figure 3 presents the normalized category co-occurrence heatmap across all models. The co-occurrence values are normalized row-wise by dividing each occurrence by the frequency of the corresponding category. Our results reveal strong inter-dependencies between inefficiency patterns in LLM-generated code. General Logic inefficiencies exhibit the highest co-occurrence rates with other categories, particularly Readability (0.74), Errors (0.64), and Maintainability (0.58). This suggests that flawed logic not only impacts correctness but also contributes to inefficiencies in clarity, structure, and long-term maintainability of the generated code. Similarly, Performance inefficiencies frequently co-occur with Readability (0.48) and Maintainability (0.46), highlighting that performance inefficiencies are often tied to code complexity and structure, rather than isolated algorithmic inefficiencies. Interestingly, Errors demonstrate weaker co-occurrences with other inefficiencies, with their strongest link being to Maintainability (0.36), indicating that while syntax and dependency errors are less common, they may still hinder maintainability efforts by requiring additional debugging and refactoring. Meanwhile, Readability and Maintainability are closely related (0.35), reinforcing the idea that clear code structures contribute to long-term maintainability. These findings suggest that improving LLM-generated logic and performance could lead to broader improvements across multiple inefficiency categories, reducing the need for manual corrections and enhancing code usability in real-world development.
@@ IV-B RQ2: Relevance of Identified Inefficiencies for Practitioners and Researchers
The survey remained open for four weeks, reaching $58$ responses, resulting in a response rate of $7.39$%, which is in line with rates observed in similar surveys in Software Engineering [39, 54, 55, 56]. First, we present the demographic information about the participants and their experience with LLM-assisted coding. Then, we provide a detailed analysis of the popularity and perceived relevance of inefficiencies in LLM-generated code reported by the participants.
@@ IV-B1 Demographic Information and Experience with Coding using LLMs
All participants answered the demographic questions. We had participants from both academia (60%) and industry (40%). The academic participants consisted of graduate students (29), undergraduate students (3), and postdoctoral researchers/lecturers (3). Industry participants included 11 software developers, 7 ML/data science developers, 4 senior executives, and 1 cybersecurity developer. About coding experience, 37 participants had more than 5 years of experience, while 11 had between 3-5 years, and 10 had between 1-3 years. LLM-assisted coding is widely adopted, with 98% of the participants using LLMs for coding. Among them, 68% rely exclusively on proprietary models, with ChatGPT and Copilot being the most popular (89.66% and 29.31%, respectively). Regarding open-source, the most popular models are Mistral, the Phi model family, and Qwen-Coder (5.17%, 3.45%, and 3.45%, respectively). Participants reported using a variety of programming languages, with the top five being Python, C/C++, JavaScript, Java, and Go (91.38%, 18.97%, 18.97%, 8.62%, and 3.45%, respectively).
Before the questions about our taxonomy, we asked participants about their general experiences and challenges with LLM-generated code. Reported challenges can be categorized into two main groups: user experience (outside the scope of our taxonomy) and code-related, which are directly aligned with our research. We discuss the inefficiencies encountered by the participants based on their experiences before seeing the taxonomy.
@@
These inefficiencies focus on the challenges/difficulties that users face when interacting with LLMs, rather than the generated code itself. 6 participants reported that LLMs struggle with debugging or modifying specific parts of code, a finding consistent with previous studies [57]. In the same way, programming language and library popularity also play a significant role in the efficiency of LLM-generated code. 5 participants noted that LLMs generate more efficient code for popular languages and libraries like Python and NumPy compared to less common ones like DHL and Tinkter. This difference is likely because LLMs are trained on more examples of popular languages and libraries, making them more familiar with their syntax and usage. Programming language paradigms also play a role—P12111We refer to survey participants using the notation Px, where x is an assigned participant ID. observed that LLMs handle interpreted languages better than compiled ones. Finally, participants expressed varying preferences for chat-based LLMs (e.g., ChatGPT) versus inline code assistants (e.g., Copilot). While P41 found Copilot suggestions as “stupid suggestions and distracted”, P11 preferred inline assistance over chat-based interactions.
@@
Participants also reported various inefficiencies in the generated code that align with the focus of our taxonomy. 19 participants stated that LLMs often struggle to produce the correct logic on the first attempt, even for simple tasks with detailed descriptions, requiring multiple further iterations. P22 specifically noted that LLMs may generate “unnecessarily long or complex code (for simple tasks)” However, P15 found that for more complex tasks such as building a Flask application, LLMs tend to perform “oversimplification to the point of uselessness”, where generation is limited to high-level steps or generalized code that lack specificity, performance efficiency, and functionality. Additionally, 3 participants reported that handling exceptions and corner cases is often overlooked. These reported inefficiencies fall under the General Logic category of our taxonomy.
Another notable challenge reported by the participants is the misalignment between LLMs’ advancements and technological progress. 13 participants reported that LLMs struggle with version tracking, frequently generating code that uses deprecated methods, mixes library versions, or is incompatible with specified APIs (also mentioned by previous works [53]). These inefficiencies align with Wrong Method Input (General Logic) and Missing Import (Errors) subcategories of our taxonomy.
Readability inefficiencies were also reported by 7 respondents spanning various programming languages, like Python, C/C++, Java, JavaScript, Go, and Bash. They highlighted poor formatting, failure to adhere to coding conventions, inconsistent variable names, and reliance on outdated practices as reported by Zhang et al. [16].
Additionally, 6 participants highlighted performance issues. Others pointed to a lack of robust, modular, object-oriented architecture and insufficient documentation and comments as well. The inefficiencies reported by the participants align with the literature [16, 14, 16] and fall under Readability, Performance, and Maintainability categories of our taxonomy.
Besides inefficiencies, 7 participants identified context limitations as a key issue. 2 participants found that LLMs had a limited context window and frequently forgot past interactions, as mentioned in related work [47]. Meanwhile, P5 reported that LLMs tended to retain outdated context, stating that they needed to “start a new chat to get rid of hallucinations.”